OpenStudy (anonymous):

Ok so can some one just like give me the equation solved like all together. I will fan and medal. I wrote it wrong last time hers the correct equation Solve the system by substitution -x-y-z=-8 -4x+4y+5z=7 2x+2z=4

jhonyy9 (jhonyy9):

do you know what mean by substitution ?

jhonyy9 (jhonyy9):

like a first step in case of 3rd equation factoriz out the 2 on both sides and divide both sides by 2 2x+2z=4 2(x+z)=4 /:2 x+z=2 x=2-z in case of first equation -x-y-z=-8 rewrite it -(x+y+z)= -(8) and divide both sides by -1 and will get x+y+z=8 so but you know that x=2-z so this you can substituting in place of x inside equation x+y+z =8 so will get 2-z+y+z=8 2+y=8 y=8-2 y=6 so you know y=6 now just you need substituting this y=6 in second equation and will get -4x+4*(6) +5z=7 -4x +24 +5z=7 -4x +5z= - 17 so but you know that x=2-z than now just you need substituting this x=2-z in place of x inside equation -4x+5z=-17 -4(2-z)+5z=-17 -8+4z+5z=-17 -8+9z=-17 9z=-9 z=-1 but you know that x=2-z so x=2-(-1) so x=2+1 so x=3 than result x=3 y=6 z=-1 hope so much that this will help you understanding all right easy good luck bye

OpenStudy (anonymous):

Thank you for your help.

jhonyy9 (jhonyy9):

np was my pleasure bye